OpenStudy (anonymous):

Becca has a biscuit recipe that uses 2/3 of a tablespoon of salt for every 3/4 of a cup of flour. In order to have enough biscuits for her family that is visiting, Becca is making a batch of biscuits using 3 cups of flour and how many tablespoons of salt will she need

OpenStudy (anonymous):

how many tablespoons of salt will that equal

hero (hero):

Okay so basically what we're given is \(\dfrac{2}{3}\) teaspoons of salt for every \(\dfrac{3}{4}\) cup of flour What you do is multiply both fractions by 4 to get: \(\dfrac{8}{3}\) teaspoons of salt for every \(3\) cups of flour

hero (hero):

Is this information clear to you?

hero (hero):

Also remember \(\dfrac{8}{3} = 2 + \dfrac{2}{3}\)